How quickly can you do this? Fill appropriate sign. ( ‘’) 
1 / 4 ☐ 2 / 8

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We know 2 / 8 = 1 / 4
Hence, 1 / 4 = 2 / 8

Thanks